# ADGaussian: Generalizable Gaussian Splatting for Autonomous Driving with Multi-modal Inputs

### v1

@@ 4 Experiments
MVSplat [6] 23.52 0.760 0.152
DepthSplat [36] 21.99 0.715 0.173
Ours 23.60 0.776 0.164
Table 2: Quantitative comparison with state of the art on KITTI dataset. Cells highlighted in denotes the best and second-best performances. Despite the inherent challenges posed by KITTI’s lower image quality and poor color reproduction, our method achieves a a small but meaningful improvement over existing approaches using only single-frame image.
4.1 Implementation Details
Datasets.
We evaluate our proposed approach on two widely used autonomous driving datasets: the Waymo Open Dataset [24] and the KITTI Tracking benchmarks [8]. For both datasets, we adopt a train-test split ratio of approximately 1:7. Specifically, on the Waymo dataset, our focus primarily lies on static and dynamic scenes, where each scene type is divided into 4 test scenes and 28 training scenes. Similarly, for the KITTI dataset, the split consists of 5 test scenes and 37 training scenes. This partitioning ensures a balanced evaluation of our method across diverse scenarios, while also providing sufficient training data for effective model training.
Metrics.
For render quality evaluation, we employ the standard image quality metrics, including Peak Signal-to-Noise Ratio (PSNR), Structural Similarity Index Measure (SSIM) [33], and the Learned Perceptual Image Patch Similarity (LPIPS) [44]. The running time and GPU memory usage are also provided to enable a comprehensive comparison of the trade-offs between speed and accuracy (see Supplementary Material for details).
Training details.
Our implementation is based on the PyTorch framework. We employ the Adam [16] optimizer and cosine learning rate schedule, with an initial learning rate of $41e-$ 4. We train our model on 3090 Ti GPUs, running for 150k iterations on both Waymo and KITTI datasets, with a batch size of 1. To ensure a fair comparison, all experiments are carried out at resolutions of 320 × 480 for the Waymo dataset and 256 × 608 for the KITTI dataset.
Figure 3: Qualitative comparison with state of the art on Waymo dataset. Our ADGaussian surpasses all other competitive models in rendering quality within urban scenarios, thanks to the efficacy of our multi-modal matching-based architecture.
4.2 Comparisons with the State of the Art
When comparing our work to the current state-of-the-art Gaussian Splatting methods, we have opted for the multi-view cost volume-based method, MVSplat [6], and the depth foundation model-based method, DepthSplat [36]. As GGRt [19] features a pose-free architecture, we excluded it from our comparisons. In each scenario, MVSplat and DepthSplat utilize two frames as input to generate subsequent novel frame for evaluation.
@@ 4.2 Comparisons with the State of the Art
The quantitative comparisons on the Waymo and KITTI benchmarks are presented in Table 1 and Table 2, respectively. On the Waymo dataset, our ADGaussian surpasses previous state-of-the-art models on almost all visual metrics, with particularly notable improvements in static scenes. Additionally, our approach demonstrates consistent performance across diverse scenarios, underscoring the robustness of its multi-modal feature matching capabilities. From Table 2, it can be observed that the performance gain of our method on the KITTI dataset is less pronounced compared to that on the Waymo dataset. This is primarily attributed to the overall lower image quality and poor color reproduction of the KITTI dataset. Since our method relies solely on a single image as input, it retains fewer image details compared to previous works, which further constrains its performance on datasets with inferior image quality.
We also provide qualitative comparisons of the two datasets in Figure 3 and Figure 4. As can be seen, our model achieves superior rendering quality, particularly in occluded regions and fine details like slender signal poles.
@@
Furthermore, the comparison between DepthSplat and MVSplat shows that DepthSplat exhibits stronger depth inference capabilities, attributed to its enhanced geometry reconstruction facilitated by pre-trained depth models. However, DepthSplat falls short in overall visual reconstruction quality due to its insufficient integration of appearance attributes, which is consistent with our earlier analysis in the preceding sections.
4.3 Ablations and Analysis
Ablations on proposed components.
The ablation studies are detailed in Table 3 to further confirm the efficacy of the proposed components. First, it can be seen that the full model achieves the highest performance, boasting a PSNR, SSIM, and LPIPS score of 31.0, 0.921, and 0.068, respectively. Notably, the removal of the depth-guided positional embedding (DPE) resulted in a decrease across all metrics (0.69, 1.3%, and 1%, respectively), emphasizing the significance of depth cues in facilitating the joint optimization of multi-modal features. Furthermore, the model lacking multi-scale Gaussian decoding (w/o Multi-scale) exhibited reduced performance, achieving a PSNR of 28.73, an SSIM of 0.868, and an LPIPS of 0.100, underscoring the effectiveness of multi-level depth decoding and independent Gaussian inference. Removing both DPE and Multi-scale led to a more substantial drop in performance, notably a 4.6% decrease in the LPIPS score.
Finally, to showcase the effectiveness of our synchronized multi-modal optimization formulation, we present the results without multi-modal feature matching (w/o Matching) by substituting the sparse depth input with a color image from the subsequent frame. It is evident that multi-modal feature matching brought about significant enhancement in PSNR, NDS, and mAP (4.32, 10.7%, and 3.8%, respectively), highlighting the importance of information exchange and synchronized optimization of image-related appearance features and depth-related geometric features.
Figure 4: Qualitative comparison with state of the art on KITTI dataset. As highlighted by the red boxes, our ADGaussian demonstrates superior performance in preserving visual consistency, particularly in handling fine details such as thin poles and object edges.
@@ 4.3 Ablations and Analysis
Table 3: Ablation studies on the Waymo dataset. We report the averaged scores across all validation scenes for a more intuitive reflection of the model performance. The “Full Model” denotes our final model and all other ablation models are based on the final model by removing the mentioned module.
@@
CFormer [45] 25.71 0.796 0.126
BPNet [27] 26.10 0.802 0.144
Ours 31.00 0.921 0.068
Table 4: Performance analysis on multi-modal inputs. Models marked with ”*” are modified with a Gaussian head for 3DGS prediction. Despite utilizing multi-modal inputs, depth completion networks still suffer from the unsatisfactory novel view rendering quality, emphasizing the importance of joint optimization and collaborative prediction strategies.
Analysis on multi-modal inputs.
To ensure a fair comparison, we constructed baseline networks with identical multi-modal inputs using state-of-the-art depth completion methods [45, 31]. Specifically, we re-implemented CFormer [45] and BPNet [27] as the comparison targets, which take both depth and image as inputs and predict Gaussian parameters using multi-modal fused features. As shown in Table 4, it is evident that the inclusion of additional depth input alone does not significantly enhance the quality of novel view rendering, further highlighting the critical need for our joint optimization. Moreover, the depth comparison results in Figure 5 reveal that our model excels in capturing fine-grained details, such as preserving the shape of cars and poles, even without pre-training on depth completion task. Additional details can be found in the Supplementary Material.
Figure 5: Depth comparison with BPNet. Our method demonstrates superior depth estimation performance in certain challenging regions, even without depth pre-training.
Cross-dataset generalization.
We perform cross-dataset evaluation to demonstrate the generalization capability of our method to out-of-distribution novel scenes. Specifically, models trained on the KITTI dataset are directly evaluated on the Waymo dataset without fine-tuning. As shown in Table 5, even with only single-view input images, our ADGaussian method consistently outperforms DepthSplat and achieves comparable cross-dataset generalization performance to MVSplat, attaining higher PSNR and SSIM scores. This result underscores our robustness in handling diverse and unseen scenarios.
@@
MVSplat [6] 17.25 0.554 0.381
DepthSplat [36] 12.47 0.441 0.497
Ours 17.63 0.557 0.432
Table 5: Cross-dataset generalization. Models trained on KITTI are directly used to test on Waymo without any further fine-tuning.
4.4 Application: Novel-view Shifting
Figure 6: Visual comparisons on view shifting. The first two rows display the performance of right and left shifting of the given images on the Waymo dataset, whereas the last two rows report visual comparisons between the Ground Truth (GT) and shifted results. The challenging areas are marked with red rectangles. As observed, our model exhibits superior robustness under large viewpoint changes.
@@ 4.4 Application: Novel-view Shifting
Table 6: Robust analysis on novel-view shifting on the KITTI dataset. Models trained on multi-frame images are utilized directly to evaluate its capability for novel-view shifting, transitioning from the left camera to the right camera, all without additional fine-tuning. Note that our ADGaussian demonstrates superior robustness compared to prior studies on novel-view shifting.
The concept of novel-view shifting involves generating images from significantly varied perspectives compared to the original viewpoints present in the training data. This task is particularly demanding as it usually necessitates reliable depth estimations to handle substantial changes in viewpoint, scale, and occlusions. In this study, we further investigate the model’s robustness in view shifting. Firstly, the ground truth right camera images provided in the KITTI dataset are used to evaluate the quantitative performance of view shifting. As depicted in Table 6, our model significantly outperforms both MVSplat and DepthSplat in zero-shot view shifting from left to right cameras. It is noteworthy that our zero-shot view shifting results are only slightly lower than our normally trained model (PSNR: 23.60, SSIM: 0.776, LPIPS: 0.164). Moreover, visual comparisons across the two datasets are presented in Figure 6. Whether shifting to the right or left, our ADGaussian exhibits exceptional overall view shifting quality on the Waymo dataset, particularly in capturing vehicle shape and intricate appearance details. Similarly, the qualitative comparisons with ground truth images on the KITTI dataset further validate the superiority of our model in view shifting.
5 Conclusion
This paper introduced a novel multi-modal framework for generalizable street scene reconstruction, demonstrating that the joint optimization of multi-modal features significantly improves both geometric and visual reconstruction quality. Additionally, we validated that our model supports zero-shot viewpoint shifting, highlighting its capabilities in generating reliable scale cues.
Limitations.
Due to its reliance on single-frame input, our method’s accuracy is limited on lower-quality datasets, and its improvements in dynamic scenes are less significant than in static scenes. These issues can be addressed in future works by effectively integrating multi-frame information.

### v2

@@ IV-B Comparisons with the State of the Art
MVSplat [2] 23.52 0.760 0.152 24.65 0.833 0.141
DepthSplat [31] 21.99 0.715 0.173 25.36 0.838 0.135
Ours 23.60 0.776 0.164 26.10 0.853 0.122
@@
The quantitative comparisons on the Waymo and KITTI benchmarks are presented in Table I and Table II, respectively. On the Waymo dataset, our ADGaussian surpasses previous state-of-the-art models on almost all visual metrics, with particularly significant gains in static scenes and consistent performance across diverse scenarios. While competing methods suffer from pervasive spatial misalignment issues that severely degrade their metric accuracy (as evidenced in Fig. 3), our model effectively resolves this fundamental challenge through cross-modal joint learning, which synchronizes geometric and appearance optimization to achieve pixel-perfect alignment. On the KITTI dataset, we introduce an additional left-to-right view synthesis setting, leveraging its stereo camera data. As shown in Table II, our method achieves superior performance in both settings compared to prior works. However, the performance gain on KITTI is less pronounced compared to that on Waymo. This is primarily attributed to the overall lower image quality and poor color reproduction of the KITTI dataset. Since our method relies solely on a single image as input, it retains fewer image details compared to previous works, which further constrains its performance on datasets with inferior image quality.
We provide qualitative results of the two datasets in Fig. 3. As can be observed, the zoomed regions in the first row reveal pervasive spatial misalignment in previous methods, which leads to significant degradation in evaluation metrics. The last two rows further demonstrate that our model achieves superior rendering quality, particularly in occluded regions and fine details such as slender signal poles.
@@
Besides, the comparison between DepthSplat and MVSplat shows that DepthSplat exhibits stronger depth inference capabilities, attributed to its enhanced geometry reconstruction facilitated by pre-trained depth models. However, DepthSplat falls short in overall visual reconstruction quality due to its insufficient integration of appearance attributes, which is consistent with our earlier analysis in the preceding sections.
IV-C Ablations and Analyses
Ablations on proposed components
TABLE III: Ablation studies on the Waymo dataset. We report the averaged scores across all validation scenes for a more intuitive reflection of model performance.
@@ IV-C Ablations and Analyses
TABLE IV: Analyses on forward view shifting and running efficiency.
@@
MVSplat [2] 24.84 0.777 0.133 0.22/0.14 11.11
DepthSplat [31] 23.40 0.726 0.190 0.37/0.28 21.17
Ours 27.68 0.877 0.101 0.29/0.18 17.52
Figure 4: Ablations on the Waymo dataset. We present a visual comparisons between our baseline model (w/o DPE & Multi-scale) and the full model. Our experiments demonstrate that the full model can effectively address the spatial misalignment issues prevalent in street scene reconstruction.
Figure 5: Depth comparisons with depth completion networks. Our method demonstrates superior depth estimation performance in certain challenging regions, even without depth pre-training.
The ablation studies are detailed in Table III to confirm the efficacy of proposed components. First, it can be seen that the full model achieves the highest performance, boasting a PSNR, SSIM, and LPIPS score of 31.0, 0.921, and 0.068, respectively. Notably, the removal of the Depth-guided Positional Embedding (DPE) resulted in a decrease across all metrics (0.69, 1.3%, and 1%, respectively), emphasizing the significance of depth positions in facilitating the joint optimization of multi-modal features. Furthermore, the model lacking Multi-scale Gaussian Decoding (w/o Multi-scale) exhibited reduced performance, achieving a PSNR of 28.73, an SSIM of 0.868, and an LPIPS of 0.100, underscoring the effectiveness of multi-level depth decoding and independent Gaussian inference. Removing both DPE and Multi-scale led to a more substantial drop in performance, notably a 4.6% decrease in the LPIPS score. This quantitative degradation aligns with the qualitative results in Fig. 4, which validates the importance of our DPE and Multi-scale Gaussian Decoding in addressing spatial misalignment for street scene reconstruction.
Finally, to showcase the effectiveness of our synchronized multi-modal optimization formulation, we present the results without Multi-modal Feature Matching (w/o Matching) by substituting the sparse depth input with a color image from the subsequent frame. It is evident that Multi-modal Feature Matching brought about significant enhancement in PSNR, SSIM, and LPIPS (4.32, 10.7%, and 3.8%, respectively), highlighting the importance of information exchange and synchronized optimization of image-related appearance features and depth-related geometric features.
Robustness across frames
In the previous experiments, the next frame (t+1 frame) was used as the target for novel view synthesis. To further evaluate the robustness under more significant viewpoint changes, we extend the prediction to the t+2 frame. As reported in Table IV, our ADGaussian outperforms in handling larger temporal and spatial shifts, even with only a single input frame.
Analysis on running efficiency
The right side of table IV shows training/inference runtime and memory consumption of ADGaussian, MVSplat and DepthSplat at 320×480 resolution to validate ADGaussian’s practicality. ADGaussian uses 3.65 GB less training memory than DepthSplat for higher efficiency, and its inference speed is merely 0.04 s slower than MVSplat with comparable overall runtime.
TABLE V: Performance analysis on multi-modal inputs. Models marked with ”*” are modified with a Gaussian head for 3DGS prediction. The term ”Waymo + Depth Drop” refers to our robustness evaluation on depth quality.
Method Waymo KITTI Waymo + Depth Drop
@@
CFormer [38] 25.71 0.796 0.126 21.35 0.761 0.192 23.67 0.780 0.141
BPNet [23] 26.10 0.802 0.144 19.68 0.626 0.336 24.89 0.783 0.152
Ours 31.00 0.921 0.068 23.60 0.776 0.164 30.56 0.912 0.074
Figure 6: Visual comparisons on view shifting. The figure displays the performance of right and left shifting of the given images on the Waymo dataset. The challenging areas are marked with red rectangles. As observed, our model exhibits superior robustness under large viewpoint changes.
IV-D Analyses on Multi-modal Inputs
To ensure a fair comparison, we further constructed baseline networks with identical multi-modal inputs using state-of-the-art depth completion methods. Specifically, we re-implemented CFormer [38] and BPNet [23] as the comparison targets, which take both image data and sparse depth as inputs and predict Gaussian parameters using multi-modal fused features. Also, we initialized these models with weights pre-trained on the KITTI depth completion dataset. By maintaining identical input modalities and training conditions, these baselines allow us to evaluate the performance gains attributable to the inclusion of depth data, independent of the architectural advancements in our framework.
Comparison Results
As shown in Table V, it is evident that the inclusion of additional depth input alone does not significantly enhance the quality of novel view rendering. This demonstrates the limitations of relying solely on accurate depth prediction for high-quality rendering, further highlighting the critical need for our joint optimization. Additionally, as demonstrated in Fig. 5, our model effectively preserves fine structural details (e.g., car contours and pole geometries) without pre-training on depth completion tasks.
Robustness to depth quality
We evaluate our model’s robustness to depth quality degradation by randomly discarding approximately 50% of LiDAR points during testing. As presented in the right part of Table V, ADGaussian remains competitive with full-depth-input performance under such conditions. The performance gap compared to baselines validates our cross-modal optimization framework’s ability to compensate for missing depth information through effective appearance-geometry integration.
IV-E Application: Novel-view Shifting
The concept of novel-view shifting involves generating images from significantly varied perspectives compared to the original viewpoints present in the training data. This task is particularly demanding as it usually necessitates reliable depth estimations to handle substantial changes in viewpoint and scale. In this study, we further investigate the model’s robustness in view shifting. Firstly, the Ground Truth right camera images provided in the KITTI dataset are used to evaluate the quantitative performance of view shifting. As depicted in Table VI, our model significantly outperforms both MVSplat and DepthSplat in zero-shot view shifting from left to right cameras. It is noteworthy that our zero-shot view shifting results are only slightly lower than our normally trained model (PSNR: 23.60, SSIM: 0.776, LPIPS: 0.164). Moreover, visual comparisons on the Waymo dataset are presented in Fig. 6. ADGaussian exhibits exceptional view shifting quality, accurately maintaining object shapes and preserving intricate details even in texture-less regions during both leftward and rightward viewpoint changes.
TABLE VI: Robust analyses on novel-view shifting on the KITTI dataset. Models trained on multi-frame images are utilized directly to evaluate their capability for novel-view shifting, transitioning from the left camera to the right camera, all without additional fine-tuning.
@@ IV-E Application: Novel-view Shifting
V Conclusion
This paper presents ADGaussian, a novel multimodal framework that advances generalizable street scene reconstruction via synergistic learning of visual and geometric features. We validate that jointly optimizing RGB imagery and sparse depth inputs significantly improves both geometric quality and visual fidelity. Comprehensive experiments on Waymo and KITTI datasets establish state-of-the-art performance in complex urban scenarios. The framework also shows strong zero-shot generalization for large viewpoint shifts while maintaining metric scale consistency. Current limitations include the sparsity of single-view information and the challenge of modeling dynamic objects. Extending the approach to multi-frame fusion could help address these gaps.